OpenStudy (anonymous):

exact value of Tan(-17pi/6)

OpenStudy (anonymous):

is it \[ \frac{1}{\sqrt{3}} \]

OpenStudy (anonymous):

\[ \tan \left(-\frac{17 \pi }{6}\right)=\tan \left(\frac{\pi }{6}-2 \pi \right)=\tan \left(+\frac{\pi }{6}\right)=\frac{1}{\sqrt{3}} \]